Find the slope and the y-intercept of each line whose equation is given
7x=5y or 7x+0=5y

Answer by stanbon(75887)

Find the slope and the y-intercept of each line whose equation is given
7x=5y or 7x+0=5y
---
These two equations are the same:
Solve for y to see the slope-intercept form:
y = (7/5)x
---
slope = 7/5
intercept = 0
===================
Cheers,
Stan H.